OpenStudy (anonymous):

Solve The Equation Using Square Roots. x(square) - 25 = 0

OpenStudy (john_es):

\[x^2=25\Rightarrow \sqrt{x^2}=\sqrt{25}\]From this point you can find the answer.

OpenStudy (anonymous):

is it 5 @John_ES

OpenStudy (anonymous):

no

OpenStudy (john_es):

No, you need the complete solution \[x=\pm 5\]It has two solutions, +5 and -5.

OpenStudy (anonymous):

how do i find the square root of x^2 lol there isnt no square root right?

OpenStudy (anonymous):

it is true that \(5^2=25\) so if you replace \(x\) by \(5\) it will work, but it will also work if \(x=-5\) as well

OpenStudy (anonymous):

@John_ES

OpenStudy (john_es):

It is easy, the square root of a square eliminates the 2 of the x.

OpenStudy (anonymous):

is that the answer or is there more to it?

OpenStudy (john_es):

I mean, if you need to find the \[\sqrt{x^2}\] you need to find the number whose square is x^2. And this is x.

OpenStudy (john_es):

No, this is the complete answer. +5 and -5.

OpenStudy (anonymous):

how is the answer suppose to look? u dont have to give me the answer just yet but i would like how the answer could look @John_ES

OpenStudy (john_es):

I would write the answer in this form \[x=\pm5\]In this concrete type of problems, the left side is always the same, and the right side is the square root of a number.

OpenStudy (anonymous):

lol thanks man i got lost because how my computer showed what u typed but i got it now its 5 is plus and minus @John_ES

OpenStudy (john_es):

Yes ;)